Jenna offers regular haircuts for $25 and haircuts plus coloring for $42. This weekend, had a total of 24 clients and Jenna earned $736.
how many regular haircuts did she do?
How many colored haircuts did she do

Answer by ankor@dixie-net.com(22740)

Jenna offers regular haircuts for $25 and haircuts plus coloring for $42.
This weekend, had a total of 24 clients and Jenna earned $736.
how many regular haircuts did she do?
How many colored haircuts did she do
:
let r = no. of regular haircuts
let c = no. of colored
:
Two equation, two unknowns, we can use substitution here
r + c = 24
r = 24 - c
and
25r + 42c = 736
replace r with (24-c)
25(24-c) + 42c = 736
600 - 25c + 42c = 736
17c = 736 - 600
17c = 136
c = 136/17
c = 8 colored haircuts
then
24 - 8 = 16 regular
:
:
See if this checks out.
16 * 25 = 400
8 * 42 = 336
----------------
total $$: 736